# On Poincaré–Sobolev level involving fractional GJMS operators on hyperbolic space

### v2

@@ 1  Introduction and Main Results
which satisfies the conformal intertwining identities (1.3)–(1.4) for all $s>0$; see [29, Theorem 1.7]. This operator provides a convenient bridge between analysis on $\mathbb{H}^{n}$ and the Euclidean setting. Moreover, the precise relation between $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ is given by [29, Corollary 5.3]:
@@ 1.1 Euclidean (Fractional) Laplacian and Integer-Order GJMS Operators
  • High dimensions relative to $s\in(0,1)$: $n\geq 4s$. One has ([33, Claim 14.1])
@@
    and consequently (1.15) admits a nontrivial weak solution for $\lambda\in(0,\lambda_{1,s}(\Omega)).$ ([33, Theorem 14.1])
  • Low dimensions relative to $s\in(0,1)$: $2s<n<4s$. There exists a constant $\lambda_{s}^{*}>0$ such that ([33, Proposition 16.4])
@@
    In this regime, the relation between $\lambda_{s}^{*}$ and $\lambda_{1,s}(\Omega)$ is unknown—even when $\Omega$ is the unit ball; consequently, the standard mountain–pass scheme alone does not guarantee existence. In [33, Theorem 16.1], the authors combine mountain–pass and linking arguments, treating separately the cases $\lambda_{s}^{*}<\lambda_{1,s}(\Omega)$ and $\lambda_{s}^{*}\geq\lambda_{1,s}(\Omega)$, and thereby obtain nontrivial solutions of (1.15) for every $\lambda>\lambda_{s}^{*}$ that is not a Dirichlet eigenvalue of $(-\Delta)^{s}$.
@@
  • If $n\geq 4$, then for $\frac{n(n-2)}{4}<\lambda\leq\frac{(n-1)^{2}}{4},$ problem (1.16) admits a positive entire solution ([30, Theorem 1.5]); for $\lambda\leq\frac{n(n-1)}{4}$, problem (1.16) does not have any positive entire solution.
@@
  • If $n\geq 4$, then $0<\mu\leq\frac{1}{4}$ guarantees a positive entire solution to (1.17) [30, Theorem 1.5]; whereas for $\mu\;\leq\;0,$ problem (1.17) admits no positive entire solution [30, Theorem 1.6].
  • If $n=3$, then for $\mu\;\leq\;\frac{1}{4},$ problem (1.17) has no positive entire solution [30, Theorem 1.7].
@@
    When $n\geq 4k$, $\mathcal{G}_{n,k}[H_{n,k}]=(0,+\infty)$. $H_{n,k}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $k\geq 2$, while for $k=1,$ it is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}]$ and strictly decreasing on $(0,\lambda_{0,k}^{\mathrm{conf}}]$, -∞=H_n,k(λ)¡0¡H_n,k(λ_0,k^conf)   for λ¿λ_0,k^conf   and   H_n,k(λ)=S_n,k for $\lambda\leq 0$.
@@ 1.2 Fractional GJMS Operators
For $s\in(0,\,n/4]\cap\mathbb{N}$, the above result recovers [28, Theorem 1.7] and yields an alternative proof thereof; whereas for $s\in(n/4,\,n/2)$, they established that $\lambda_{s}^{\mathrm{conf}}\in\bigl(0,\,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr).$
@@
    For $s\in\left(0,\frac{n}{4}\right]$, $\mathcal{G}_{n,s}[\widetilde{H}_{n,s}]=(0,+\infty)$, $\widetilde{H}_{n,s}(\lambda)$ is achieved and strictly decreasing for $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})$, -∞=~H_n,s(μ)¡0¡~H_n,s(~λ_0,s^conf)   for μ¿~λ_0,s^conf   and   ~H_n,s(λ)=S_n,s for $\lambda\leq 0$.
@@
We conjecture that, in the regime $s\in(0,\frac{n}{4}]$, the level $\widetilde{H}_{n,s}(\lambda)$ is achieved if and only if $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$, and when $n\geq 3$, $s\in[\frac{n-1}{2},\frac{n}{2}),$ $\widetilde{H}_{n,s}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$.
@@
    For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{+}$, $(b_{s},\infty)\subset\mathcal{G}_{n,s}\bigl[H_{n,s}\bigr]$ and $H_{n,s}(\lambda)$ is achieved and strictly decreasing for $\lambda\in(b_{s},\lambda_{0,s}^{\mathrm{conf}})$. Moreover, H_n,s(λ)=S_n,s    for $\lambda\leq 0$.
@@
(iii) When $n\geq 3$ and $s\in[\frac{n-1}{2},\frac{n}{2}),$ we conjecture that problem (1.9) admits no positive solution for any $\lambda\in\mathbb{R}$, thus $\widetilde{H}_{n,s}$ is never achieved; for $s\in(0,\frac{n}{4}]$, we conjecture that $\widetilde{H}_{n,s}(\lambda)$ is achieved if and only if $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$.
@@
    (Subcritical case) If $1<p<2_{s}^{*}-1$, assume that $\lambda<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$.
@@
    $\mathcal{G}_{n,s}\!\big[\widetilde{H}_{n,s}\big]\cap(-\infty,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})\neq\varnothing\qquad\text{and}\qquad\lambda\in\mathcal{G}_{n,s}\!\big[\widetilde{H}_{n,s}\big]\cap(-\infty,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}).$
@@
    $\mathcal{G}_{n,s}\!\big[\widetilde{H}_{n,s}\big]\cap(-\infty,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})=(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}).$
@@
    (Subcritical case) If $1<p<2_{s}^{*}-1$, assume that $\lambda<\lambda_{0,s}^{\mathrm{conf}}$.
@@
    $\mathcal{G}_{n,s}\!\big[H_{n,s}\big]\cap(-\infty,\lambda_{0,s}^{\mathrm{conf}})\neq\varnothing\qquad\text{and}\qquad\lambda\in\mathcal{G}_{n,s}\!\big[H_{n,s}\big]\cap(-\infty,\lambda_{0,s}^{\mathrm{conf}}).$
@@
    $(b_{s},\lambda_{0,s}^{\mathrm{conf}})\subset\mathcal{G}_{n,s}\!\big[H_{n,s}\big]\cap(-\infty,\lambda_{0,s}^{\mathrm{conf}}).$
@@ 2.2 Helgason Fourier Transform
which coincides with Sobolev space $H^{2}(\mathbb{B}^{n})$, see [23, Appendix B].
@@
The spectral theorem for multiplication operators on $L^{2}$ implies that the spectrum of $M_{\varphi}$ is the essential range of its symbol $\varphi$, see [23, Appendix A]. Since $\mathcal{H}$ is unitary, the spectrum is invariant under this unitary equivalence,
@@ 2.3 Fractional GJMS Operators and Inequality
However, $\mathcal{P}_{\gamma}$ does not satisfy the following intertwining relations (1.3) and (1.4). To recover a direct intertwining with the Euclidean fractional Laplacian under the conformal identifications of $\mathbb{H}^{n}$ with $\mathbb{R}^{n}_{+}$ and $\mathbb{B}^{n}$, Lu et al. introduced an auxiliary family of fractional operators $\widetilde{\mathcal{P}}_{\gamma}$ on $\mathbb{H}^{n}$ in [29, Theorem 1.7]. Equivalently, under the Helgason–Fourier transform one has the multiplier representation
@@ 3.1 Asymptotic Analysis
Combining this estimate with [29, Theorem 1.8], we can only conclude that
@@
(i) Applying [29, Lemma 5.2] with $\gamma=s$ and $\lambda=0$, we obtain
@@ 4.1 Energy Spaces
and by [23, Theorem E.8], we obtain the sharp fractional Poincaré inequality
@@
By [29, Theorem 1.3 and Theorem 1.8], for every $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s}$ defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$, and for every $\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s,\sim}$ also defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$
@@
$\|u\|_{\lambda,s}:=\sqrt{\langle(\mathcal{P}_{s}-\lambda)u,u\rangle_{L^{2}(\mathbb{H}^{n})}},\quad\forall u\in C_{c}^{\infty}(\mathbb{H}^{n})$
@@
Set $\lambda_{0}:=\lambda_{0,s}^{\mathrm{conf}}$ and $\delta:=\lambda_{0}-\lambda>0$. Since $m_{s}(\beta)\geq\lambda_{0}$ for all $\beta$, we have $m_{s}(\beta)-\lambda\geq\delta$. Moreover,
@@
We define the energy space as the completion of $C_{c}^{\infty}(\mathbb{H}^{n})$ under the norm $\|\cdot\|_{\lambda,s}$, namely
@@
$\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}):=\overline{C_{c}^{\infty}(\mathbb{H}^{n})}^{\ \|\cdot\|_{\lambda,s}}\quad{\rm for}\ \,\lambda<\lambda_{0,s}^{\mathrm{conf}}.$
@@
We study the attainability of fractional Sobolev levels associated with $P_{s}-\lambda$ and $\widetilde{P}_{s}-\lambda$
@@
$H_{n,s,p}(\lambda):=\inf_{u\in\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})\setminus\{0\}}\frac{\mathcal{E}_{\lambda,s}(u)}{\|u\|_{L^{p+1}(\mathbb{H}^{n})}^{2}}\in(0,\infty),\quad\lambda<\lambda_{0,s}^{\mathrm{conf}},$ (4.6)
@@
$\widetilde{H}_{n,s,p}(\lambda):=\inf_{u\in\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})\setminus\{0\}}\frac{{\widetilde{\mathcal{E}}}_{\lambda,s}(u)}{\|u\|_{L^{p+1}(\mathbb{H}^{n})}^{2}}\in(0,\infty),\quad\lambda<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}.$
For convenience, we set

### v3

@@ 1  Introduction and Main Results
which satisfies the conformal intertwining identities (1.3)–(1.4) for all $s>0$; see [30, Theorem 1.7]. This operator provides a convenient bridge between analysis on $\mathbb{H}^{n}$ and the Euclidean setting. Moreover, the precise relation between $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ is given by [30, Corollary 5.3]:
@@ 1.1 Euclidean (Fractional) Laplacian and Integer-Order GJMS Operators
  • High dimensions relative to $s\in(0,1)$: $n\geq 4s$. One has ([34, Claim 14.1])
@@
    and consequently (1.15) admits a nontrivial weak solution for $\lambda\in(0,\lambda_{1,s}(\Omega)).$ ([34, Theorem 14.1])
  • Low dimensions relative to $s\in(0,1)$: $2s<n<4s$. There exists a constant $\lambda_{s}^{*}>0$ such that ([34, Proposition 16.4])
@@
    In this regime, the relation between $\lambda_{s}^{*}$ and $\lambda_{1,s}(\Omega)$ is unknown—even when $\Omega$ is the unit ball; consequently, the standard mountain–pass scheme alone does not guarantee existence. In [34, Theorem 16.1], the authors combine mountain–pass and linking arguments, treating separately the cases $\lambda_{s}^{*}<\lambda_{1,s}(\Omega)$ and $\lambda_{s}^{*}\geq\lambda_{1,s}(\Omega)$, and thereby obtain nontrivial solutions of (1.15) for every $\lambda>\lambda_{s}^{*}$ that is not a Dirichlet eigenvalue of $(-\Delta)^{s}$.
@@
  • If $n\geq 4$, then for $\frac{n(n-2)}{4}<\lambda\leq\frac{(n-1)^{2}}{4},$ problem (1.16) admits a positive entire solution ([31, Theorem 1.5]); for $\lambda\leq\frac{n(n-1)}{4}$, problem (1.16) does not have any positive entire solution.
@@
  • If $n\geq 4$, then $0<\mu\leq\frac{1}{4}$ guarantees a positive entire solution to (1.17) [31, Theorem 1.5]; whereas for $\mu\;\leq\;0,$ problem (1.17) admits no positive entire solution [31, Theorem 1.6].
  • If $n=3$, then for $\mu\;\leq\;\frac{1}{4},$ problem (1.17) has no positive entire solution [31, Theorem 1.7].
@@
    When $n\geq 4k$, $\mathcal{G}_{n,k}[H_{n,k}]=(0,+\infty)$. $H_{n,k}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $k\geq 2$; while for $k=1,$ it is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}]$ and strictly decreasing on $(0,\lambda_{0,k}^{\mathrm{conf}}]$, -∞=H_n,k(λ)¡0¡H_n,k(λ_0,k^conf)   for λ¿λ_0,k^conf   and   H_n,k(λ)=S_n,k for $\lambda\leq 0$.
@@ 1.2 Fractional GJMS Operators
For $s\in(0,\,n/4]\cap\mathbb{N}$, the above result recovers [29, Theorem 1.7] and yields an alternative proof thereof; whereas for $s\in(n/4,\,n/2)$, they established that $\lambda_{s}^{\mathrm{conf}}\in\bigl(0,\,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr).$
@@
    For $s\in\left(0,\frac{n}{4}\right]$, $\mathcal{G}_{n,s}[\widetilde{H}_{n,s}]=(0,+\infty)$, $\widetilde{H}_{n,s}(\lambda)$ is achieved if and only if $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$ and strictly decreasing in $(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$. Moreover, -∞=~H_n,s(μ)¡0¡~H_n,s(~λ_0,s^conf)   for μ¿~λ_0,s^conf   and   ~H_n,s(λ)=S_n,s for $\lambda\leq 0$.
@@
We conjecture that, when $n\geq 3$, $s\in[\frac{n-1}{2},\frac{n}{2}),$ $\widetilde{H}_{n,s}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$.
@@
    For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{+}$, $(b_{s},\infty)\subset\mathcal{G}_{n,s}\bigl[H_{n,s}\bigr]$ and $H_{n,s}(\lambda)$ is achieved and strictly decreasing for $\lambda\in(b_{s},\lambda_{0,s}^{\mathrm{conf}}]$. Moreover, H_n,s(λ)=S_n,s    for $\lambda\leq 0$.
@@
(iii) When $n\geq 3$ and $s\in[\frac{n-1}{2},\frac{n}{2}),$ we conjecture that problem (1.9) admits no positive solution for any $\lambda\in\mathbb{R}$, thus $\widetilde{H}_{n,s}$ is never achieved.
@@
    (Subcritical case) If $1<p<2_{s}^{*}-1$, assume that $\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$.
@@
    $\mathcal{G}_{n,s}\!\big[\widetilde{H}_{n,s}\big]\cap(-\infty,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]\neq\varnothing\qquad\text{and}\qquad\lambda\in\mathcal{G}_{n,s}\!\big[\widetilde{H}_{n,s}\big]\cap(-\infty,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}].$
@@
    $\mathcal{G}_{n,s}\!\big[\widetilde{H}_{n,s}\big]\cap(-\infty,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]=(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}].$
@@
    (Subcritical case) If $1<p<2_{s}^{*}-1$, assume that $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$.
@@
    $\mathcal{G}_{n,s}\!\big[H_{n,s}\big]\cap(-\infty,\lambda_{0,s}^{\mathrm{conf}}]\neq\varnothing\qquad\text{and}\qquad\lambda\in\mathcal{G}_{n,s}\!\big[H_{n,s}\big]\cap(-\infty,\lambda_{0,s}^{\mathrm{conf}}].$
@@
    $(b_{s},\lambda_{0,s}^{\mathrm{conf}}]\subset\mathcal{G}_{n,s}\!\big[H_{n,s}\big]\cap(-\infty,\lambda_{0,s}^{\mathrm{conf}}].$
@@ 2.2 Helgason Fourier Transform
which coincides with Sobolev space $H^{2}(\mathbb{B}^{n})$, see [24, Appendix B].
@@
The spectral theorem for multiplication operators on $L^{2}$ implies that the spectrum of $M_{\varphi}$ is the essential range of its symbol $\varphi$, see [24, Appendix A]. Since $\mathcal{H}$ is unitary, the spectrum is invariant under this unitary equivalence,
@@ 2.3 Fractional GJMS Operators and Inequality
However, $\mathcal{P}_{\gamma}$ does not satisfy the following intertwining relations (1.3) and (1.4). To recover a direct intertwining with the Euclidean fractional Laplacian under the conformal identifications of $\mathbb{H}^{n}$ with $\mathbb{R}^{n}_{+}$ and $\mathbb{B}^{n}$, Lu et al. introduced an auxiliary family of fractional operators $\widetilde{\mathcal{P}}_{\gamma}$ on $\mathbb{H}^{n}$ in [30, Theorem 1.7]. Equivalently, under the Helgason–Fourier transform one has the multiplier representation
@@ 3.1 Asymptotic Analysis
Combining this estimate with [30, Theorem 1.8], we can only conclude that
@@
(i) Applying [30, Lemma 5.2] with $\gamma=s$ and $\lambda=0$, we obtain
@@ 4.1 Energy Spaces
and by [24, Theorem E.8], we obtain the sharp fractional Poincaré inequality
@@
By [30, Theorem 1.3 and Theorem 1.8], for every $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s}$ defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$, and for every $\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s,\sim}$ also defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$
@@
$\|u\|_{\lambda,s}=\sqrt{\langle(\mathcal{P}_{s}-\lambda)u,u\rangle_{L^{2}(\mathbb{H}^{n})}},\quad\forall u\in C_{c}^{\infty}(\mathbb{H}^{n})$
@@
Set $\lambda_{0}:=\lambda_{0,s}^{\mathrm{conf}}$ and $\delta:=\lambda_{0}-\lambda>0$. Since $m_{s}(\beta)\geq\lambda_{0}\geq 0$ for all $\beta$, we have $m_{s}(\beta)-\lambda\geq\delta$. Moreover,
@@
For $s>0$ and $\lambda\in\mathbb{R}$, we define the energy spaces $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$ and $\mathcal{H}^{s}_{\lambda,\sim}(\mathbb{H}^{n})$ as the completions of $C_{c}^{\infty}(\mathbb{H}^{n})$ with respect to the norms $\|\cdot\|_{\lambda,s}$ and $\|\cdot\|_{\lambda,s,\sim}$, respectively, i.e.
$\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}):=\overline{C_{c}^{\infty}(\mathbb{H}^{n})}^{\,\|\cdot\|_{\lambda,s}},\qquad\mathcal{H}^{s}_{\lambda,\sim}(\mathbb{H}^{n}):=\overline{C_{c}^{\infty}(\mathbb{H}^{n})}^{\,\|\cdot\|_{\lambda,s,\sim}}.$
Remark 4.1.
Let $\lambda_{0,s}^{\mathrm{conf}}$ and $\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$ denote the bottom spectral values associated with $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$, respectively. By Lemma 4.2, if $\lambda<\lambda_{0,s}^{\mathrm{conf}}$, then
$\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})=H^{s}(\mathbb{H}^{n}).$
By Lemma 4.1, if $\lambda<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, then
$\mathcal{H}^{s}_{\lambda,\sim}(\mathbb{H}^{n})=H^{s}(\mathbb{H}^{n}).$
In both cases, the corresponding energy norm is equivalent to the standard $H^{s}(\mathbb{H}^{n})$ norm.
Lemma 4.4.
Let $\lambda$ be the corresponding bottom spectral value, namely
@@
$\lambda=\lambda_{0,s}^{\mathrm{conf}}\quad\text{for }\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}),\qquad\text{or}\qquad\lambda=\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\quad\text{for }\mathcal{H}^{s}_{\lambda,\sim}(\mathbb{H}^{n}).$
Then the following statements hold.
(i) If
$u\in\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})\quad\text{(respectively, }u\in\mathcal{H}^{s}_{\lambda,\sim}(\mathbb{H}^{n})\text{)}$
and $\operatorname{supp}u\Subset\mathbb{H}^{n}$, then $u\in H^{s}(\mathbb{H}^{n}).$
(ii) Let $\{v_{n}\}\subset\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$ (respectively, $\{v_{n}\}\subset\mathcal{H}^{s}_{\lambda,\sim}(\mathbb{H}^{n})$) satisfy
$v_{n}\rightharpoonup 0\quad\text{weakly in }\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}),\quad\textnormal{(respectively, weakly in $\mathcal{H}^{s}_{\lambda,\sim}(\mathbb{H}^{n})$).}$
Assume moreover that there exists a compact set $K\Subset\mathbb{H}^{n}$ such that
$\operatorname{supp}v_{n}\subset K,\qquad\forall n.$
Then
$v_{n}\rightharpoonup 0\quad\text{weakly in }H^{s}(\mathbb{H}^{n}).$
Proof: We only prove the statement for $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$ with $\lambda=\lambda_{0,s}^{\mathrm{conf}}$, since the proof for $\mathcal{H}^{s}_{\lambda,\sim}(\mathbb{H}^{n})$ with $\lambda=\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$ is identical (replacing $\mathcal{P}_{s}$ by $\widetilde{\mathcal{P}}_{s}$ throughout).
(i) Let $u\in\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$ with $\operatorname{supp}u\Subset\mathbb{H}^{n}$. By definition, there exists a sequence $\{u_{m}\}\subset C_{c}^{\infty}(\mathbb{H}^{n})$ such that
$u_{m}\to u\quad\text{in }\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}).$
Since $\operatorname{supp}u\Subset\mathbb{H}^{n}$, we may choose a compact set $K\Subset\mathbb{H}^{n}$ with $\operatorname{supp}u\subset K$, and we may assume
$\operatorname{supp}u_{m}\subset K,\qquad\forall m.$
By [30, Theorem 1.3], there exists $q>2$ and $C>0$ such that for all $w\in C_{c}^{\infty}(\mathbb{H}^{n})$,
$\|w\|_{L^{q}(\mathbb{H}^{n})}\leq C\,\|w\|_{\lambda,s}.$
Applying this to $w=u_{m}-u_{\ell}$, and using $u_{m}\to u$ in $\mathcal{H}^{s}_{\lambda}$, we obtain that $\{u_{m}\}$ is Cauchy in $L^{q}(\mathbb{H}^{n})$. Since all $u_{m}$ are supported in the fixed compact set $K$, Hölder’s inequality yields
$\|u_{m}-u_{\ell}\|_{L^{2}(\mathbb{H}^{n})}=\|u_{m}-u_{\ell}\|_{L^{2}(K)}\leq|K|^{\frac{1}{2}-\frac{1}{q}}\|u_{m}-u_{\ell}\|_{L^{q}(K)}\to 0.$
Hence $\{u_{m}\}$ is Cauchy in $L^{2}(\mathbb{H}^{n})$.
We now compare the critical energy norm with the $H^{s}$-norm. Let $m_{s}(\beta)$ denote the spectral multiplier of $\mathcal{P}_{s}$; then $m_{s}(0)=\lambda_{0,s}^{\mathrm{conf}}=\lambda$, and by (4.4) one has
$m_{s}(\beta)-\lambda\sim(\beta^{2}+\rho^{2})^{s}\quad\text{for large }|\beta|,$
while adding the $L^{2}$ term restores the low frequencies. Equivalently, there exists $C_{1},C_{2}>0$ such that
$C_{1}\Bigl(1+(\beta^{2}+\rho^{2})^{s}\Bigr)\leq 1+\bigl(m_{s}(\beta)-\lambda\bigr)\leq C_{2}\Bigl(1+(\beta^{2}+\rho^{2})^{s}\Bigr),\qquad\forall\beta\in\mathbb{R}.$
Therefore, for every $w\in C_{c}^{\infty}(\mathbb{H}^{n})$,
$\|w\|_{H^{s}(\mathbb{H}^{n})}^{2}\sim\|w\|_{L^{2}(\mathbb{H}^{n})}^{2}+\|w\|_{\lambda,s}^{2}.$
Applying this to $w=u_{m}-u_{\ell}$ and using that $\{u_{m}\}$ is Cauchy both in $L^{2}$ and in $\mathcal{H}^{s}_{\lambda}$, we conclude that $\{u_{m}\}$ is Cauchy in $H^{s}(\mathbb{H}^{n})$. Hence $u_{m}\to\tilde{u}$ in $H^{s}(\mathbb{H}^{n})$ for some $\tilde{u}\in H^{s}(\mathbb{H}^{n})$. By [30, Theorem 1.3] and [5, Theorem 6.1], we obtain the limit is unique, so $\tilde{u}=u$. Therefore, $u\in H^{s}(\mathbb{H}^{n}).$
(ii) Assume $v_{n}\rightharpoonup 0$ weakly in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$ and $\operatorname{supp}v_{n}\subset K\Subset\mathbb{H}^{n}$ for all $n$. Weak convergence implies boundedness, so
$\sup_{n}\|v_{n}\|_{\lambda,s}<\infty.$
By [30, Theorem 1.3] and the common compact support, there exists $q>2$ such that
$\sup_{n}\|v_{n}\|_{L^{q}(\mathbb{H}^{n})}\leq C\sup_{n}\|v_{n}\|_{\lambda,s}<\infty,$
and hence
$\sup_{n}\|v_{n}\|_{L^{2}(\mathbb{H}^{n})}=\sup_{n}\|v_{n}\|_{L^{2}(K)}\leq|K|^{\frac{1}{2}-\frac{1}{q}}\sup_{n}\|v_{n}\|_{L^{q}(K)}<\infty.$
Using again
[…]
we infer that $\{v_{n}\}$ is bounded in $H^{s}(\mathbb{H}^{n})$.
To prove the weak convergence of the whole sequence in $H^{s}(\mathbb{H}^{n})$, it suffices to show that every subsequence admits a further subsequence converging weakly to 0 in $H^{s}(\mathbb{H}^{n})$.
Let $\{v_{n_{j}}\}$ be an arbitrary subsequence of $\{v_{n}\}$. Since $\{v_{n}\}$ is bounded in $H^{s}(\mathbb{H}^{n})$, so is $\{v_{n_{j}}\}$; hence, there exist a further subsequence (still denoted by $\{v_{n_{j}}\}$) and some $w\in H^{s}(\mathbb{H}^{n})$ such that
$v_{n_{j}}\rightharpoonup w\qquad\text{weakly in }H^{s}(\mathbb{H}^{n}).$
In particular, $v_{n_{j}}\to w$ in $\mathcal{D}^{\prime}(\mathbb{H}^{n})$. On the other hand, since $v_{n}\rightharpoonup 0$ in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, we also have $v_{n_{j}}\to 0$ in $\mathcal{D}^{\prime}(\mathbb{H}^{n})$. Therefore $w=0$. As the chosen subsequence $\{v_{n_{j}}\}$ was arbitrary, every subsequence of $\{v_{n}\}$ admits a further subsequence converging weakly to 0 in $H^{s}(\mathbb{H}^{n})$. Hence
$v_{n}\rightharpoonup 0\qquad\text{weakly in }H^{s}(\mathbb{H}^{n}).$
This proves (ii). ∎
@@
Next, we begin to study the attainability of fractional Sobolev levels associated with $\mathcal{P}_{s}-\lambda$ and $\widetilde{\mathcal{P}}_{s}-\lambda$
@@
$H_{n,s,p}(\lambda):=\inf_{u\in\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})\setminus\{0\}}\frac{\mathcal{E}_{\lambda,s}(u)}{\|u\|_{L^{p+1}(\mathbb{H}^{n})}^{2}}\in(0,\infty),\quad\lambda\leq\lambda_{0,s}^{\mathrm{conf}},$ (4.6)
@@
$\widetilde{H}_{n,s,p}(\lambda):=\inf_{u\in\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})\setminus\{0\}}\frac{{\widetilde{\mathcal{E}}}_{\lambda,s}(u)}{\|u\|_{L^{p+1}(\mathbb{H}^{n})}^{2}}\in(0,\infty),\quad\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}.$
In the critical case $p+1=2_{s}^{*}:=\frac{2n}{n-2s}$, we note that
$H_{n,s,p}(\lambda)=H_{n,s}(\lambda)\qquad\text{and}\qquad\widetilde{H}_{n,s,p}(\lambda)=\widetilde{H}_{n,s}(\lambda),$
where $H_{n,s}(\lambda)$ and $\widetilde{H}_{n,s}(\lambda)$ are defined in (1.1) and (1.7). For convenience, we set